# Description length of canonical and microcanonical models

### v2

@@ IV Applications to maximum-entropy models
In the case of bipartite binary networks, the sequence $r_{i}(\mathbf{G})$, $i=1\dots n$ coincides with the degree sequence of the nodes belonging to just one layer and induces the canonical MEM known as Bipartite Partial Configuration Model saracco17 . As we will employ asymptotic results to compare the canonical and the microcanonical description lengths, we need to specify the scaling of the constraints $\mathbf{c}^{*}$ as $n\to\infty$. In what follows, we will focus on dense matrices, i.e. we will assume that $g_{ij}^{*}=\Theta(1)$ $\forall(i,j)$, where $\Theta(x)$ indicates a quantity of the same order of $x$ as $n\to\infty$. This choice leads to $r_{i}^{*}=\Theta(m)$ and $l^{*}=\Theta(nm)$, the ‘actual’ order of these quantities depending on the growth rate of $m$. We will consider two cases: (i) $m$ finite in the limit, i.e. $m=\Theta(1)$ and (ii) $m$ growing linearly with $n$, i.e., $m=\Theta(n)$. As a consequence of our assumption, the matrix density
@@
will remain finite in the asymptotic limit, i.e. $p^{*}=\Theta(1)$ and $p_{i}^{*}=\Theta(1)$, $\forall\>i$.
@@ IV.1 Enforcing one global constraint
$P_{\text{mic}}(\mathbf{G};l)=\begin{dcases}\frac{1}{\binom{nm}{l}}&\text{if}\>l(\mathbf{G})=l\\ 0&\text{else},\end{dcases}$ (42)
@@
i.e. the distribution of a collection of $nm$ i.i.d. Bernoulli variables, indicating that each matrix element is either $1$, with probability $p$, or 00, with probability $1-p$. The ML estimator of $p$ is given by the matrix density $\hat{p}(\mathbf{G})=l(\mathbf{G})/nm$ and the maximum log-likelihood, obtained by inserting $\hat{p}$ into (48), reads
@@
where $h(p)=-p\ln(p)-(1-p)\ln(1-p)$ is the entropy of a Bernoulli distribution with a parameter equal to $p$. An exact expression for the canonical complexity has been derived in staniczenko and reads
@@
So far, all the reported results are exact. Nevertheless, as we are interested in the order of $\Delta\text{DL}$ when $n\to\infty$, we consider the following asymptotic results, derived from the exact ones (see A.1):
@@
The expressions above provide two significant insights. First, the order of the difference between log-likelihood functions is $o(n)$, regardless of the growth rate of $m$, a result implying that condition (36) is met and ensemble equivalence holds. Second, the leading terms of expressions (53) and (54) are identical, thus cancelling out when compared. Consequently, the (asymptotic) difference between description lengths depends on the size of the system only through the matrix density $p^{*}$, assumed to be finite in the dense regime. Thus, in the same regime, such a difference is finite, in agreement with (37).
@@ IV.2 Enforcing $n$ local constraints
$P_{\text{mic}}(\mathbf{G};\mathbf{r})=\begin{dcases}\prod_{i=1}^{n}\frac{1}{\binom{m}{r_{i}}}&\text{if}\>\mathbf{r}(\mathbf{G})=\mathbf{r}\\ 0&\text{else}\end{dcases}$ (56)
@@
In the hypothesis of $m$ growing linearly with $n$, i.e. $m=\Theta(n)$, we consider the following asymptotic expressions (see A.2)
@@
As for the case of one global constraint, we focus on two aspects of these results. First, the difference between log-likelihood functions is either $\Theta(n)$ or $\Theta(n\ln n)$ depending on the order of $m$, i.e. $m=\Theta(1)$ in the first case and $m=\Theta(n)$ in the second case. In any case, according to (36), the ensemble equivalence is broken. In particular, if $m=\Theta(n)$, the leading terms of expressions (IV.2) and (IV.2) are both of order $\Theta(n\ln n)$, thus cancelling out when compared - precisely as in the equivalent case. Nonetheless, the remaining terms are still of order $\Theta(n)$. Hence, the (asymptotic) difference between description lengths increases linearly with the system size, regardless of the growth rate of $m$.
@@ IV.3 Future research: Towards model-level equivalence
In this section, we explicitly show that, in both cases considered, the difference between likelihoods and the difference between complexities share the exact same leading order. This confirms that the contribution of the complexities in comparing the two formulations is as important as that of the likelihoods.
Furthermore, relation (37) is confirmed in the case of one global constraint examined here, where equivalence holds and the difference between the canonical and microcanonical description lengths turns out to be $o(n)$. At the same time, the result that $\Delta\text{DL}$ is $\Theta(n)$ in the case of $n$ local constraints shows that the signature of the breaking of ensemble equivalence is still visible when the models’ complexities are involved in the comparison.
These findings offer valuable insights for future research on non-equivalence between microcanonical and canonical MEMs within the model selection framework. They lead to the following question: can we generalize the definition of equivalence at the level of the models, i.e. by including their complexities and independently of the observed data?
$\Delta\text{DL}$ is not suitable for this purpose: while it does take into account the models’ complexities, it still depends on the observed values. Instead, we propose a definition involving the universal NML distributions representing the two models. Inspired by definition 35, we consider the following KL divergence
@@
This relation states that whenever ensemble equivalence holds, the corresponding microcanonical and canonical models are equivalent on the model level. The reverse implication, verified here in the specific case of one-sided local constraints, will be investigated in future work.
@@ V NML-based VS Bayesian approach to model selection
In the previous sections, we implicitly assume that no prior knowledge about the parameters is available. Under this assumption, the NML-based universal distribution is the best choice in a precise minimax sense, as shown in section II. Here, we consider the Bayesian approach to model selection grunwald_book ; grunwald2019 by introducing the Bayesian universal distribution
@@
(see  peixoto17 ; vallescatala18 for applications in the context of network models). The main difference with respect to the NML description length lies in the presence of a prior on the parameters $w(\bm{\theta})$ to be supplied by the user. This implementation, which is formally equivalent to the Bayes factor method kass95 , provides a description length for $\bm{x}$ through model $\mathcal{M}$ that is defined as
@@
where $\text{det}\mathbf{I}(\theta)$ is the determinant of the $k\times k$ normalized Fisher information matrix. If we equip the Bayesian universal distribution with the Jeffreys prior jeffreys1946
@@
by posing $w(\bm{\theta})\equiv J(\bm{\theta})$, the two asymptotic expressions coincide (for a formal derivation of these results, we redirect the interested reader to  rissanen96 and  grunwald_book ). Thus, whenever the asymptotic formulas (V) and (V) hold true, the NML-based description length is asymptotically equivalent to the one provided by the Bayesian distribution equipped with Jeffreys prior (hereby, Bayes-Jeffreys description), which is NML-optimal.
@@
(see B.1). In the limit $n\to\infty$ one finds
@@
As $n\to\infty$, the last difference diverges both in case $m$ stays finite, i.e. $m=\Theta(1)$, and in case $m$ diverges, upon assuming that the rate at which $m$ diverges does not exceed that of $n$. Consequently, in this case, it is no longer true that the NML approach can be asymptotically retrieved from the Bayesian one equipped with Jeffreys prior. Although the existence of a modified Jeffreys prior extending the equivalence of the two approaches to the case of $n$ local constraints can indeed be hypothesized, for the time being, we merely suggest not to assume this to be true when ensemble non-equivalence holds.
In the remainder of this section, we show that the Bayesian description length becomes more sensitive to the prior choice whenever local constraints are enforced. Let us start by considering the relationship
@@
stating that the NML-based description length of a canonical model coincides with the Bayesian description length of its microcanonical variant, equipped with the so-called canonical prior, introduced by Rissanen in rissanen2001 and defined in B.2. Moreover,
@@
i.e. the NML-based description length of the microcanonical model coincides with the Bayesian description length of its canonical variant, equipped with a uniform prior on the mean-value parameters.
The two identities, derived in B.2, show that there exists a prior guaranteeing that the Bayesian description length of the canonical (microcanonical) model coincides with the NML-based description length of the microcanonical (canonical) model. This might lead to the conclusion that, in the framework of model selection, canonical and microcanonical models coincide even when non-equivalence holds if the right priors are chosen. Nevertheless, these priors are not NML-optimal. While this non-optimality is negligible in the case of one global constraint, it could play a crucial role when local constraints are enforced. In other words, in this case, different choices of prior can result in very different Bayesian description lengths for microcanonical as for canonical models.
@@
This simple example shows that when non-equivalence holds, the selection of different priors can result in increasingly larger differences in DL when the system size grows.
@@ VI Conclusion
This study examined the implications of ensemble non-equivalence within MDL model selection. Specifically, we compared the NML-based description lengths of canonical and microcanonical maximum-entropy models.
@@
We found that while microcanonical models fit the data better than their canonical counterparts, they are also more complex. The model achieving the minimum description length depends on the observed values of the sufficient statistics.
@@
We explicitly compared the description lengths of canonical and microcanonical models of dense $n\times m$ binary matrices in a couple of cases of interest: one global constraint, such as the sum of the entries of the matrix, and $n$ local constraints, such as the sum of the entries of each row of the matrix. While the former induces a homogeneous model, according to which the probability of an entry being 1 is the same for all entries, the latter induces a heterogeneous model, according to which the probability of an entry being 1 depends on the specific row. In the second scenario, microcanonical and canonical ensembles are no longer equivalent. By comparing the description length of the canonical and the microcanonical formulations of these two models, we observed that their difference remains finite when one global constraint is enforced while growing linearly with the system size when $n$ local constraints are enforced.
@@
Our findings shed light on the consequences that may arise from choosing either hard or soft constraints in terms of data compression, leading us to conclude that canonical and microcanonical models should be regarded as distinct from a model selection perspective in the presence of ensemble non-equivalence. Moreover, they suggest a possible new definition of model equivalence based on MDL that needs further investigation.
@@
Furthermore, we explored the relationship between NML-based and Bayesian description lengths. While the Jeffreys prior is NML-optimal for canonical models under a single global constraint, this is no longer true when local constraints are enforced. In this case, the difference between these two description lengths diverges as the system size increases. This observation should warn practitioners regarding the two approaches as interchangeable in the presence of ensemble non-equivalence, and it raises the question of whether an NML-optimal prior exists when local constraints are enforced.
@@
Lastly, we showed that different priors can result in largely different description lengths when local constraints are involved. This outcome underscores the importance of carefully selecting priors when dealing with an extensive number of constraints.
Appendix A Asymptotic expansions
@@
Here, we derive the asymptotic expansions used to obtain the asymptotic results in section IV.
A.1 One global constraint
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-\ln\binom{nm}{l}=-\ln(nm)!+\ln(nm-l)!+\ln l!.$ (87)
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-nm\ln(nm)+(nm-l)\ln(nm-l)+l\ln l-\frac{1}{2}\ln\left(\frac{nm}{2\pi l(nm-l)}\right)+o(1),$ (89)
@@
$\ln\mathcal{L}_{\text{mic}}(p)=-nm\cdot h(p)+\frac{1}{2}\ln(nm)+\frac{1}{2}\ln(2\pi p(1-p))+o(1).$ (90)
@@
$\ln\mathcal{L}_{\text{can}}(p)=-nm\cdot h(p),$ (91)
@@
$\text{COMP}_{\text{mic}}$ $=\ln(nm+1)=\ln(nm)+\ln\left(1+\frac{1}{nm}\right)$ (92)
$=\ln(nm)+o(1),$
@@
$\text{COMP}_{\text{can}}=\frac{1}{2}\ln(nm)+\frac{1}{2}\ln\left(\frac{\pi}{2}\right)+o(1).$ (93)
@@
$\text{COMP}_{\textrm{can}}^{\text{NML}}\simeq\ln\left[\sqrt{\frac{\pi nm}{2}}+\frac{2}{3}+\frac{\sqrt{2\pi}}{24\sqrt{nm}}-\frac{4}{135nm}+\frac{\sqrt{2\pi}}{576\sqrt{(nm)^{3}}}+\frac{8}{2835(nm)^{2}}\right],$ (94)
provided in  [25], where the authors show it to be very precise already for small values of the total number of entries $nm$. Finally, equation (54) is obtained by comparing the two asymptotic complexities. Notice that the asymptotic formula (93) for the canonical complexity represents a very well-known result, as it can be derived directly by applying the asymptotic formula (V) to the case of i.i.d. Bernoulli random variables.
A.2 One-sided local constraints
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-\sum_{i=1}^{n}\ln\binom{m}{r_{i}}=-n\ln m!+\sum_{i=1}^{n}\ln(m-r_{i})!+\ln r_{i}!.$ (95)
@@
$\ln x!=x\ln x-x+\frac{1}{2}\ln(2\pi x)+\frac{1}{12x}+\Theta(1/x^{3}).$ (96)
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-m\sum_{i=1}^{n}h(p_{i})+\frac{n}{2}\ln m+\frac{1}{2}\sum_{i=1}^{n}[\ln(2\pi p_{i}(1-p_{i}))]\\ -\frac{n}{12m}+\frac{1}{12m}\sum_{i=1}^{n}\left[\frac{1}{p_{i}(1-p_{i})}\right]+o(1).$ (97)
@@
$\ln\mathcal{L}_{\text{can}}(\mathbf{p})=-m\sum_{i=1}^{n}h(p_{i}),$ (98)
@@
$=n\ln m+\frac{n}{m}+o(1),$ (99)
which follows from $\ln\left(1+\frac{1}{x}\right)=\frac{1}{x}+\Theta(1/x^{2})$, is compared to the following asymptotic expansion of the canonical complexity, based on (94)
$\text{COMP}_{\text{can}}=\frac{n}{2}\ln m+\frac{n}{2}\ln\frac{\pi}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi}}\cdot\frac{n}{\sqrt{m}}+\left(\frac{1}{12}-\frac{4}{9\pi}\right)\cdot\frac{n}{m}+o(1)$ (100)
@@
Appendix B NML and Bayes
B.1 Bayesian-Jeffreys description lengths
In what follows, we compute the Bayesian description length $\text{DL}^{\text{B, Jeffreys}}_{\text{can}}$ of the canonical model obtained by constraining the sum $l$ over the matrix. As already stated, this model is equivalent to modelling $nm$ i.i.d. Bernoulli variable, and this result can be found in Example 8.3 of [18]. Later, we extend this result to the case of $n$ one-sided local constraints.
First, we compute Jeffreys prior according to (80). For a Bernoulli variable, the Fisher information for the parameter $p$ is $\mathbf{I}(p)=p^{-1}(1-p)^{-1}$ and $\int_{0}^{1}\sqrt{\mathbf{I}(p)}=\pi$. Thus, the Jeffreys prior reads
$J(p)=\frac{1}{\pi\sqrt{p(1-p)}}$ (101)
and, according to definition (75), the Bayesian-Jeffreys DL is
@@ B.1 Bayesian-Jeffreys description lengths
$\text{DL}^{\text{B, Jeff}}_{\text{can}}(\mathbf{G}^{*})=-\ln\frac{\Gamma\left(l^{*}+\frac{1}{2}\right)\Gamma\left(nm-l^{*}+\frac{1}{2}\right)}{\pi(nm)!},$ (105)

### v3

@@ IV Applications to maximum-entropy models
In the case of bipartite binary networks, the sequence $r_{i}(\mathbf{G})$, $i=1\dots n$ coincides with the degree sequence of the nodes belonging to just one layer and induces the canonical MEM known as Bipartite Partial Configuration Model [24]. As we will employ asymptotic results to compare the canonical and the microcanonical description lengths, we need to specify the scaling of the constraints $\mathbf{c}^{*}$ as $n\to\infty$. In what follows, we will focus on dense matrices, i.e.,we will assume that $g_{ij}^{*}=\Theta(1)$ $\forall(i,j)$, where $\Theta(x)$ indicates a quantity of the same order of $x$ as $n\to\infty$. This choice leads to $r_{i}^{*}=\Theta(m)$ and $l^{*}=\Theta(nm)$, the ‘actual’ order of these quantities depending on the growth rate of $m$. We will consider two cases: (i) $m$ finite in the limit, i.e.,$m=\Theta(1)$ and (ii) $m$ growing linearly with $n$, i.e., $m=\Theta(n)$. As a consequence of our assumption, the matrix density
@@
will remain finite in the asymptotic limit, i.e.,$p^{*}=\Theta(1)$ and $p_{i}^{*}=\Theta(1)$, $\forall\>i$.
@@ IV.1 Enforcing one global constraint
$P_{\text{mic}}(\mathbf{G};l)=\begin{cases}\frac{1}{\binom{nm}{l}}&\text{if}\>l(\mathbf{G})=l\\ 0&\text{else},\end{cases}$ (42)
@@
i.e.,the distribution of a collection of $nm$ i.i.d. Bernoulli variables, indicating that each matrix element is either $1$, with probability $p$, or 00, with probability $1-p$. The ML estimator of $p$ is given by the matrix density $\hat{p}(\mathbf{G})=l(\mathbf{G})/nm$ and the maximum log-likelihood, obtained by inserting $\hat{p}$ into (48), reads
@@
where $h(p)=-p\ln(p)-(1-p)\ln(1-p)$ is the entropy of a Bernoulli distribution with a parameter equal to $p$. An exact expression for the canonical complexity has been derived in [25] and reads
@@
So far, all the reported results are exact. Nevertheless, as we are interested in the order of $\Delta\text{DL}$ when $n\to\infty$, we consider the following asymptotic results, derived from the exact ones (see section S1.1 of the supplementary material):
@@
The expressions above provide two significant insights. First, the order of the difference between log-likelihood functions is $o(n)$, regardless of the growth rate of $m$, a result implying that condition (36) is met and ensemble equivalence holds. Second, the leading terms of expressions (53) and (54) are identical, thus cancelling out when compared. Consequently, the (asymptotic) difference between description lengths depends on the size of the system only through the matrix density $p^{*}$, assumed to be finite in the dense regime. Thus, such a difference is finite in the same regime, in agreement with (37).
@@ IV.2 Enforcing $n$ local constraints
$P_{\text{mic}}(\mathbf{G};\mathbf{r})=\begin{cases}\prod_{i=1}^{n}\frac{1}{\binom{m}{r_{i}}}&\text{if}\>\mathbf{r}(\mathbf{G})=\mathbf{r}\\ 0&\text{else}\end{cases}$ (56)
@@
In the hypothesis of $m$ growing linearly with $n$, i.e.,$m=\Theta(n)$, we consider the following asymptotic expressions (for a derivation, see section S1.2 of the supplementary material)
@@
As for the case of one global constraint, we focus on two aspects of these results. First, the difference between log-likelihood functions is either $\Theta(n)$ or $\Theta(n\ln n)$ depending on the order of $m$, i.e.,$m=\Theta(1)$ in the first case and $m=\Theta(n)$ in the second case. In any case, according to (36), the ensemble equivalence is broken. In particular, if $m=\Theta(n)$, the leading terms of expressions (IV.2) and (IV.2) are both of order $\Theta(n\ln n)$, thus cancelling out when compared - precisely as in the equivalent case. Nonetheless, the remaining terms are still of order $\Theta(n)$. Hence, the (asymptotic) difference between description lengths increases linearly with the system size, regardless of the growth rate of $m$.
@@ IV.3 Future research: Towards model-level equivalence
At this point, it is important to emphasize that comparing canonical and microcanonical ensembles is not the same as comparing the correspondent models. The former involves comparing two probability distributions, each obtained for specific parameter values, whereas the latter involves comparing two sets of distributions (i.e., two models), with each set sharing a common functional form within itself. While, in the context of the measure-level definition of ensemble equivalence, the comparison between ensembles can be reduced to comparing their likelihood terms, the comparison between models must take into account both their likelihoods and complexities.
In this section, we have shown two examples where the leading-order terms of the differences in likelihood and complexity are identical. This confirms that the complexities play as crucial a role in model comparison as the likelihoods. Furthermore, we validated relation (37) in the case of a single global constraint: when ensemble equivalence holds, the difference in description lengths between the canonical and microcanonical ensembles turns out to be $o(n)$. At the same time, our result that $\Delta\text{DL}$ grows with $n$ in the case of $n$ local constraints shows that the signature of broken ensemble equivalence remains visible when model complexities are included in the comparison.
These findings lead to the following question: can we generalize the definition of ensemble equivalence at the model level? To do so, we need a quantity that plays a role analogous to that of the relative entropy when comparing ensembles but which applies to the comparison of models.
$\Delta\text{DL}$ might seem like a good candidate but is not suitable for this purpose: while it does take into account the model complexities, it still depends on the observed values. To be more general, we propose a definition involving the universal NML distributions representing the two models. Inspired by the measure level definition of non equivalence [6] of equation (35), we consider the following KL divergence
@@
i.e.,whenever ensemble equivalence holds, the corresponding microcanonical and canonical models are equivalent on the model level. The reverse implication, verified here in the specific case of one-sided local constraints, will be investigated in future work.
@@ V NML-based VS Bayesian approach to model selection
In the previous sections, we implicitly assumed that no prior knowledge about the parameters was available. Under this assumption, the NML-based universal distribution is the best choice in a precise minimax sense, as shown in section II. Here, we consider the Bayesian approach to model selection [18, 19] by introducing the Bayesian universal distribution
@@
(see  [26, 27] for applications in the context of network models). The main difference with respect to the NML description length lies in the presence of a prior on the parameters $w(\bm{\theta})$ to be supplied by the user. This implementation, which is formally equivalent to the Bayes factor method [28], provides a description length for $\bm{x}$ through model $\mathcal{M}$ that is defined as
@@
where $\text{det}\mathbf{I}(\theta)$ is the determinant of the $k\times k$ normalized Fisher information matrix.
If we equip the Bayesian universal distribution with the Jeffreys prior [29]
@@
by posing $w(\bm{\theta})\equiv J(\bm{\theta})$, the two asymptotic expressions coincide (for a formal derivation of these results, we redirect the interested reader to  [21] and  [18]). Thus, whenever the asymptotic formulas (V) and (V) hold true, the NML-based description length is asymptotically equivalent to the one provided by the Bayesian distribution equipped with Jeffreys prior (hereby, Bayes-Jeffreys description), which is NML-optimal.
@@
(see section S2.1 of the supplementary material for a derivation). In the limit $n\to\infty$ one finds
@@
As $n\to\infty$, the last difference diverges both in case $m$ stays finite, i.e.,$m=\Theta(1)$, and in case $m$ diverges, upon assuming that the rate at which $m$ diverges does not exceed that of $n$. Consequently, in this case, it is no longer true that the NML approach can be asymptotically retrieved from the Bayesian one equipped with Jeffreys prior. Although the existence of a modified Jeffreys prior extending the equivalence of the two approaches to the case of $n$ local constraints can indeed be hypothesized, for the time being, we merely suggest not to assume this to be true when ensemble non-equivalence holds.
In the remainder of this section, we show that the Bayesian description length becomes more sensitive to the choice of prior whenever local constraints are enforced. Let us start by considering the relationship
@@
stating that the NML-based description length of a canonical model coincides with the Bayesian description length of its microcanonical variant, equipped with the so-called canonical prior, introduced by Rissanen in [30] (see the definition in section S2.2 of the supplementary material). Moreover,
@@
i.e., the NML-based description length of the microcanonical model coincides with the Bayesian description length of its canonical variant equipped with a uniform prior on the mean-value parameters.
The two identities, derived in S2.2, show that there exists a prior guaranteeing that the Bayesian description length of the canonical (microcanonical) model coincides with the NML-based description length of the microcanonical (canonical) model. This might lead to the conclusion that, in the framework of model selection, canonical and microcanonical models coincide even when non-equivalence holds if the right priors are chosen. Nevertheless, these priors are not NML-optimal. While this non-optimality is negligible in the case of one global constraint, it could play a crucial role when local constraints are enforced. In other words, in this case, different choices of prior can result in very different Bayesian description lengths for microcanonical as for canonical models.
@@
This simple example shows that when non-equivalence holds, the selection of different priors can result in increasingly larger differences in DL when the system size grows. In other words, fine-tuning the prior can result in a significant gain - or, conversely, a loss - in terms of compression, particularly for large-scale systems.
@@ VI Conclusion
This study explored the implications of ensemble non-equivalence in the context of MDL-based model selection. We specifically compared the NML-based description lengths of canonical and microcanonical maximum-entropy models. The NML approach enables a comparison of model description lengths derived from the same underlying principle without requiring prior assumptions about the parameters.
@@
We found that, although microcanonical models fit the data better than their canonical counterparts, they are also more complex. Determining the best-scoring model - i.e., the one that minimizes the description length - is not straightforward, as it depends on the observed values of the sufficient statistics. Specifically, we found that the canonical model performs best when its fit to the observed data exceeds its average fit across all possible values of the sufficient statistics.
@@
Additionally, we explicitly compared the description lengths of canonical and microcanonical models for dense $n\times m$ binary matrices in two cases of interest: one with a global constraint, such as the sum of the matrix entries, and another with $n$ local constraints, such as the sum of the entries in each row. In the former case, the model is homogeneous, where the probability of an entry being 1 is the same for all entries, while in the latter, the model becomes heterogeneous, with the probability of an entry being 1 depending on the specific row. In this second scenario, the canonical and microcanonical ensembles are no longer equivalent.
@@
By comparing the description lengths of the canonical and microcanonical formulations of these models, we found that the difference remains finite when a global constraint is enforced but grows linearly with system size when local constraints are imposed. Our findings highlight the impact of choosing hard or soft constraints on data compression, leading us to conclude that canonical and microcanonical models should be treated as distinct from a model selection perspective in the presence of ensemble non-equivalence. Furthermore, they suggest a new definition of model equivalence based on MDL, which requires further investigation.
@@
In addition, we investigated the relationship between NML-based and Bayesian description lengths. We defined NML-optimal priors as those that, when applied in a Bayesian framework, result in description lengths that asymptotically match the NML-based approach. While the Jeffreys prior is well known to be NML-optimal for canonical models under a single global constraint, we show that this optimality does not hold when local constraints are imposed. Indeed, in such cases, the difference between the corresponding description lengths increases with system size. This observation serves as a caution to practitioners, suggesting that the two approaches (NML and Jeffreys priors) should not be treated as interchangeable in the presence of ensemble non-equivalence. It also raises the question of whether an NML-optimal prior exists when local constraints are enforced.
@@
Lastly, we showed that different priors can result in largely different description lengths when local constraints are involved. This finding underscores the importance for practitioners of carefully selecting priors, since the choice of prior could greatly impact the accuracy and effectiveness of model selection.
Acknowledgements
This work is supported by the European Union - NextGenerationEU - National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR), project ‘SoBigData.it - Strengthening the Italian RI for Social Mining and Big Data Analytics’ - Grant IR0000013 (n. 3264, 28/12/2021) (https://pnrr.sobigdata.it/).
Supplementary Materials
S1 Asymptotic Expansions
@@ Acknowledgements
Here we derive the asymptotic expansions used to obtain the asymptotic results in section IV.
S1.1 One global constraint
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-\ln\binom{nm}{l}=-\ln(nm)!+\ln(nm-l)!+\ln l!.$ (S1)
@@
$\ln\mathcal{L}_{\text{mic}}(l)=-nm\ln(nm)+(nm-l)\ln(nm-l)+l\ln l-\frac{1}{2}\ln\left(\frac{nm}{2\pi l(nm-l)}\right)+o(1),$ (S3)
@@
$\ln\mathcal{L}_{\text{mic}}(p)=-nm\cdot h(p)+\frac{1}{2}\ln(nm)+\frac{1}{2}\ln(2\pi p(1-p))+o(1).$ (S4)
@@
$\ln\mathcal{L}_{\text{can}}(p)=-nm\cdot h(p),$ (S5)
@@
$\text{COMP}_{\text{mic}}$ $=\ln(nm+1)=\ln(nm)+\ln\left(1+\frac{1}{nm}\right)$
$=\ln(nm)+o(1),$ (S6)
@@
$\text{COMP}_{\text{can}}=\frac{1}{2}\ln(nm)+\frac{1}{2}\ln\left(\frac{\pi}{2}\right)+o(1).$ (S7)
@@
$\text{COMP}_{\textrm{can}}^{\text{NML}}\simeq\ln\left[\sqrt{\frac{\pi nm}{2}}+\frac{2}{3}+\frac{\sqrt{2\pi}}{24\sqrt{nm}}-\frac{4}{135nm}+\frac{\sqrt{2\pi}}{576\sqrt{(nm)^{3}}}+\frac{8}{2835(nm)^{2}}\right],$ (S8)
provided in  [25], where the authors show it to be very precise already for small values of the total number of entries $nm$. Finally, equation (54) is obtained by comparing the two asymptotic complexities. Notice that the asymptotic formula (S7) for the canonical complexity represents a well-known result, as it can be derived directly by applying the asymptotic formula (V) to the case of i.i.d. Bernoulli random variables.
S1.2 One-sided local constraints
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-\sum_{i=1}^{n}\ln\binom{m}{r_{i}}=-n\ln m!+\sum_{i=1}^{n}\ln(m-r_{i})!+\ln r_{i}!.$ (S9)
@@
$\ln x!=x\ln x-x+\frac{1}{2}\ln(2\pi x)+\frac{1}{12x}+\Theta(1/x^{3}).$ (S10)
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-m\sum_{i=1}^{n}h(p_{i})+\frac{n}{2}\ln m+\frac{1}{2}\sum_{i=1}^{n}[\ln(2\pi p_{i}(1-p_{i}))]\\ -\frac{n}{12m}+\frac{1}{12m}\sum_{i=1}^{n}\left[\frac{1}{p_{i}(1-p_{i})}\right]+o(1).$ (S11)
@@
$\ln\mathcal{L}_{\text{can}}(\mathbf{p})=-m\sum_{i=1}^{n}h(p_{i}),$ (S12)
@@
$=n\ln m+\frac{n}{m}+o(1),$ (S13)
which follows from $\ln\left(1+\frac{1}{x}\right)=\frac{1}{x}+\Theta(1/x^{2})$, is compared to the following asymptotic expansion of the canonical complexity, based on (S8)
$\text{COMP}_{\text{can}}=\frac{n}{2}\ln m+\frac{n}{2}\ln\frac{\pi}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi}}\cdot\frac{n}{\sqrt{m}}+\left(\frac{1}{12}-\frac{4}{9\pi}\right)\cdot\frac{n}{m}+o(1)$ (S14)
@@
S2 NML and Bayes
S2.1 Bayesian-Jeffreys description lengths
In what follows, we compute the Bayesian description length $\text{DL}^{\text{B, Jeffreys}}_{\text{can}}$ of the canonical model obtained by constraining the sum $l$ over the matrix. As already stated, this model is equivalent to modelling $nm$ i.i.d. Bernoulli variable, and this result can be found in Example 8.3 of [18]. In the paper, we extend this result to the case of $n$ one-sided local constraints.
First, we compute Jeffreys prior. For a Bernoulli variable, the Fisher information for the parameter $p$ is $\mathbf{I}(p)=p^{-1}(1-p)^{-1}$ and $\int_{0}^{1}\sqrt{\mathbf{I}(p)}=\pi$. Thus, the Jeffreys prior reads
$J(p)=\frac{1}{\pi\sqrt{p(1-p)}}$ (S15)
and the Bayesian-Jeffreys DL is
@@
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=-\ln\frac{\Gamma\left(l^{*}+\frac{1}{2}\right)\Gamma\left(nm-l^{*}+\frac{1}{2}\right)}{\pi(nm)!},$ (S19)